Determine the equation of the circle with radius 9

9 and center (-1,-8)

(−1,−8).



1
Expert's answer
2022-03-31T02:18:25-0400

"(x-a)^2 +(y - b)^2 = R^2\\"

So we have:

"(x+1)^2+(y+8)^2 = 81"